11123:

A box has 5 beads of the same size, but all are different colors. Tina draws a bead randomly from the box, notes its color, and then puts the bead back in the box. She repeats this 3 times. What is the probability that Tina would pick a red bead on the first draw, then a green bead, and finally a red bead again? 1 over 625 1 over 180 1 over 150 1 over 125

AnimeGhoul8863:

srry didnt see the post when u want someone to answer ur question tag them like this @ name and they will appear but make sure the @ and name are together

AnimeGhoul8863:

1/125

AnimeGhoul8863:

this is the answer hope this helps !

mikewwe13:

probability of picking red on first draw = 1/5 but then the bead is put back probability of picking green = 1/5 but then the bead is put back probability of picking red again = 1/5 1/5 * 1/5 * 1/5 = 1/125